### v1

# PersonaCraft: Personalized Full-Body Image Synthesis for Multiple Identities from Single References Using 3D-Model-Conditioned Diffusion

Abstract
Personalized image generation has been significantly advanced, enabling the creation of highly realistic and customized images. However, existing methods often struggle with generating images of multiple people due to occlusions and fail to accurately personalize full-body shapes. In this paper, we propose PersonaCraft, a novel approach that combines diffusion models with 3D human modeling to address these limitations. Our method effectively manages occlusions by incorporating 3D-aware pose conditioning with SMPLx-ControlNet and accurately personalizes human full-body shapes through SMPLx fitting. Additionally, PersonaCraft enables user-defined body shape adjustments, adding flexibility for individual body customization. Experimental results demonstrate the superior performance of PersonaCraft in generating high-quality, realistic images of multiple individuals while resolving occlusion issues, thus establishing a new standard for multi-person personalized image synthesis. Project page: https://gwang-kim.github.io/persona_craft
Figure 1: PersonaCraft generates realistic, personalized images of multiple people with complex occlusions, preserving full body details, including facial identity and body shape from single references. It also supports a novel application: user-defined body shape control, allowing custom body shape adjustments.
11footnotetext: Authors contributed equally.  Corresponding author.
1 Introduction
Recent advances in personalized image generation [20, 31, 33, 60, 61, 65, 17, 4, 23, 69, 70, 56, 82, 83, 10, 18, 25, 52, 64, 85, 74, 24, 71, 48, 78, 19, 84, 64, 80] have enabled the creation of visually compelling and highly customized portraits. Diffusion models [34, 66, 84, 62, 14, 42, 39, 41, 40] have been especially impactful, significantly enhancing the adaptability and quality of generative methods through improved controllability, such as pose-guided generation with ControlNet [81]. However, generating images with multiple individuals remains challenging in spite of its development [11, 30, 50, 26, 46, 28, 58, 75, 43], particularly in complex scenarios with occlusions, where overlapping body parts or interacting poses obscure person-specific details (yellow boxes in the baselines in Fig. 1(a)).
Full-body personalization also poses significant challenges, as most existing methods primarily focus on facial identity, often overlooking individual body shape details. Many approaches rely on 2D pose estimations [79, 16, 29], which lack the depth required to capture unique physical attributes accurately. Furthermore, current methods struggle to effectively separate a person’s identity from their clothing, resulting in unwanted style transfers and inconsistencies, particularly during occluded interactions or when depicting intricate garments.
To address these limitations, we propose PersonaCraft, a framework for generating highly personalized, full-body images of multiple individuals in complex scenes. PersonaCraft integrates diffusion models with 3D human modeling (SMPLx model [57]) for precise control over body shape and pose. The approach includes three core mechanisms: 1) Face and body identity extraction: We extract body shape parameters via SMPLx fitting [13, 21, 2] and obtain face ID embeddings from reference images using InsightFace [1]. 2) 3D-aware pose conditioning with SMPLx-ControlNet (SCNet): SCNet, which leverages SMPLx depth maps and is trained on our curated MPII-SMPLx dataset featuring complex, occluded multi-human scenes, enables robust control over body shape and pose, allowing realistic interactions between individuals. 3) Multiple Human Personalized Image Synthesis: For the personalization of multiple identities, PersonaCraft uses seamlessly SCNet and IdentityNet [71], leveraging face masks to accurately localize identities. Unlike two-stage methods (e.g., OMG [43]) or test-time optimization approaches (e.g., DreamBooth [60]), PersonaCraft efficiently personalizes multiple identities within a single generative process.
PersonaCraft generates highly realistic and personalized images of people, even in occluded scenarios. As shown in Fig. 1(a), prior methods struggle with body shape personalization and anatomical inaccuracies. In contrast, PersonaCraft ensures accurate body shapes and face identities while maintaining distinct styles.
Moreover, PersonaCraft introduces a unique feature: User-defined body shape control, allowing users to adjust attributes like height and body build for tailored body shapes as shown in Fig. 1(b). This feature is valuable for applications in advertising and social media, where customization boosts engagement.
Extensive quantitative and qualitative evaluations, including user studies, validate PersonaCraft’s superior performance in preserving facial identity, body shape, and natural human anatomy, even in challenging scenarios involving multiple people and significant occlusions. These capabilities position PersonaCraft as a state-of-the-art solution for high-quality, multi-person image generation, surpassing existing methods in accuracy and personalization.
Figure 2: Overview of PersonaCraft. We extract face and body embeddings from reference images, then perform personalized image synthesis using controllable modules for face identity and body shape customization. Our method applies 3D-aware pose conditioning with SMPLx-ControlNet (SCNet) to generate occlusion-robust images that accurately reflect body shapes.
2 Related Works
Single-concept personalization. Text-to-image (T2I) diffusion models and applications [34, 66, 84, 62, 14, 42, 39, 41, 40] enable single-concept personalization, adapting pre-trained models for individual subjects. Early methods used optimization-based techniques [20, 31, 33, 60, 61, 65, 17] or textual embeddings [4, 23, 69, 70, 56, 82, 83]. LoRA methods [35, 68] reduced the need for many trainable parameters. Recent works[10, 18, 25, 52, 64, 85, 74, 24, 71, 48, 78, 19, 84, 64, 80] like IP-Adapter [80] and InstantID [71] adopts modular design enabling fast personalization from a single reference image and included human-centric modules. However, these methods often struggle with full-body fidelity, either neglecting body shape or failing to disentangle identity from clothing in the case of single reference personalization.
Multi-concept personalization. Current approaches often combine multi-concept datasets through joint training, adding losses or optimization steps to facilitate model merging. Some methods [11, 30, 50, 26] use cross-attention maps to avoid entangling concepts. Custom Diffusion [46] proposes either joint training or constrained optimization for model integration. Mix-of-show [28] utilize gradient fusion reduces identity loss, while regionally controllable sampling manages attribute binding. Modular Customization [58] leverages orthogonal directions to isolate concepts for easier integration. FastComposer [75] enhances text conditioning in diffusion models with subject embeddings, allowing personalization without retraining. OMG [43] proposes the method without additional training for multi-concept personalization enhancing efficiency and flexibility. However, existing methods still exhibit limitations in maintaining occlusion robustness and precise control over body shapes.
Pose-guidance for image synthesis. Recent works such as ControlNet [81] and HyperHuman [49] combine T2I diffusion with 2D skeleton poses for controllable human image synthesis. DWpose [79] improves on OpenPose [16] by providing more accurate skeletons to enhance diffusion models’ quality [22, 36], while DensePose [29] maps correspondences between RGB images and 3D surfaces. PODIA-3D [41] uses depth maps for pose-guidnace but not for the human image synthesis. The SMPL [51] and SMPLx [57] model , widely used for realistic human body and shape modeling, provide detailed 3D pose and body shape parameters. Champ [86] employs SMPL as a 3D condition for video synthesis but is limited to single-human cases.
Unlike previous methods, our approach uniquely handles full-body personalization with occlusion robustness and 3D body shape conditioning. Using SMPLx ControlNet (SCNet), we synthesize realistic full-body images that adapt to complex poses and occlusions, preserving face and body identity. Our method integrates seamlessly with single-reference images, enabling efficient personalization. Additionally, it offers user control over body shapes, creating a versatile human-centric generation framework that addresses limitations in existing T2I personalization techniques.
3 Proposed Method: PersonaCraft
Figure 3: SMPLx-ControlNet (SCNet) leverages 3D-aware pose conditioning using SMPLx models to accurately capture body shape and pose. The SMPLx depth map, generated from body shape parameters and 3D poses, serves as a precise conditioning signal for robust image synthesis, effectively handling occlusions.
We aim to generate high-quality, personalized images of multiple people in a complex scene from the single full-body references. Given a set of full-body reference images $\{I_{\text{ref}}^{(i)}\}^{N_{\text{human}}}_{i=1}$, target poses $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$, where $N_{\text{human}}$ is the number of human identities, and a text prompt $y$, our goal is to synthesize a personalized image $I_{P}$ that accurately reflects the identities, poses, and the semantic content specified by the text prompt. This task is particularly challenging due to complex occlusions and the need for precise full-body shape personalization.
Our proposed method, PersonaCraft, tackles these challenges using a 3D human model-based diffusion model, as shown in Fig. 2. We begin by extracting face and body identity embeddings from the reference images. Next, we apply a 3D-aware pose conditioning technique with SMPLx-ControlNet (SCNet) to generate high-quality, occlusion-robust images. We then perform personalized image synthesis for multiple humans. Finally, we introduce a novel application of our method: user-defined body shape control, allowing users to specify desired body shapes and generate customized human images.
3.1 Face and Body Identity Extraction
We first extract face and body identity information from a single reference image $I_{\text{ref}}^{(i)}$ per individual as illustrated in the left part of Fig. 2. We employ MultiHMR [13], a state-of-the-art SMPLx [57] fitting method, to perform robust 3D human reconstruction, estimating SMPLx parameters from a single image. This process extracts body shape parameters $\bm{\beta}^{(i)}$ for each person, ensuring accurate shape representation. For face identity, we utilize InstantID [71]’s face ID embedding method via InsightFace [1], obtaining face embeddings $\bm{f}^{(i)}$. These combined embeddings, $\bm{\beta}^{(i)}$ and $\bm{f}^{(i)}$, capture person-specific identity features, enabling the generation of highly personalized multi-human images.
Figure 4: Qualitative comparison of our method with the baselines demonstrates PersonaCraft’s superior performance in multi-identity images with occlusions. Our approach maintains consistent identities and accurate body shapes across varying poses, outperforming the baseline in handling occlusions, preserving body shape, and ensuring pose fidelity, even as the number of subjects increases.
3.2 3D-Aware Pose Conditioning using SMPLx-ControlNet
Existing human personalization methods often utilize 2D skeleton-based pose ControlNets [81], for pose-guided image generation. However, 2D pose estimation lacks the depth and specificity required to accurately capture unique body contours and structures, often leading to a loss of individual physical attributes. This limitation hinders full human personalization, including both face identity and body shape, and makes the models vulnerable to occlusions.
To address these issues, we leverage 3D pose information, $\bm{p}\in{N_{\text{joint}}\times 3}$, along with camera view pose $\bm{c}$, where $N_{\text{joint}}$ is the number of joints. 3D pose provides a more comprehensive representation of human pose, capturing depth information which can be significant in scenarios with complex poses and occlusions. However, even with 3D pose, naively using 3D skeleton-based pose conditioning may not fully capture body shape variations.
To overcome this limitation, we propose a novel 3D-aware pose conditioning technique using SMPLx-ControlNet (SCNet). By leveraging 3D human models, specifically SMPLx [57], we can accurately represent body shape and pose. Given body shape parameters $\{\bm{\beta}^{(i)}\}^{N_{\text{human}}}_{i=1}$, 3D poses $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$, and a camera view pose $\bm{c}$, we first generate a 3D human model using SMPLx. We then render the SMPLx depth map $\bm{d}_{\text{SMPLx}}$, which serves as a strong conditioning signal for the diffusion model and effectively handles occlusions. This SMPLx depth map is used as a precise guidance for image generation with SCNet.
MPII-SMPLx Dataset. We train SCNet by fine-tuning depth ControlNet [81], originally trained on a large text-image-depth triplet dataset, using our curated MPII-SMPLx dataset. This dataset builds upon the MPII dataset [9], a comprehensive collection of human activity images featuring diverse and challenging poses and complex interactions. Notably, there is currently no well-curated multi-human image dataset with 3D human reconstruction, making our MPII-SMPLx dataset one of our key contribution. To enhance the MPII dataset, we incorporate textual descriptions from Khan et al.[38]. For 3D human reconstruction, we employ MultiHMR[13], a state-of-the-art multi-human SMPLx [57] fitting method, to handle multiple identities per image and render depth maps from the reconstructed SMPLx meshes. To account for occlusion scenarios, we balance the dataset with a 1:2 ratio of single-person to multi-person images. Depth clipping is applied during depth rendering to retain only values below 5, ensuring consistent quality. After preprocessing, we curate a final dataset of 6,348 image-text-SMPLx-parameter pairs. This carefully curated dataset enables robust model training with diverse 3D human poses, complex interactions, and detailed human parameters such as body shape and pose conditioning.
Training objective of SCNet. Following the ControlNet [81] framework, the training objective minimizes the difference between generated and ground truth images. Given an input image $\bm{x}_{0}$, its noisy counterpart $\bm{x}_{t}$ from noise $\epsilon$ at diffusion timestep $t$, a text prompt $y$, condition $\bm{c}$, and control signal $\bm{d}_{\text{SMPLx}}$, the SCNet $\mathcal{E}^{\text{SC}}_{\phi}$ is fine-tuned from depth ControlNet using the objective:
$\vspace{-3pt}\mathbb{E}_{\bm{x}_{0},t,\bm{c},\bm{d}_{\text{SMPLx}},\epsilon}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{x}_{t},{t},\bm{c},\mathcal{E}^{\text{SC}}_{\phi}({\bm{d}_{\text{SMPLx}}}))\|_{2}^{2}\Big{]},$ (1)
where $\epsilon_{\theta}$ is a pretrained diffusion model, kept frozen during SCNet training, and $\epsilon\sim\mathcal{N}(0,1)$.
Figure 5: Additional comparison of generated images. Yellow arrows highlight anatomical issues due to 2D pose limitations. Blue arrows refer to the individuals evaluated for correct body shape preservation. PersonaCraft excels in identity, body shape consistency, and naturalness while being over twice as fast as OMG-based methods [71, 80, 43].
3.3 Multiple Human Personalized Image Synthesis
To synthesize personalized images of multiple individuals, we leverage a face ControlNet, IdentitiyNet  [72], and SCNet. To ensure precise identity preservation and enhanced image quality, we employ face masks to accurately localize facial regions from the given pose. Suppose we aim to generate a personalized image, given a set of $N_{\text{human}}$ reference images $\{I_{\text{ref}}^{(i)}\}^{N_{\text{human}}}_{i=1}$, corresponding target poses $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$, and a text prompt $y$. Then, we obtain face embeddings $\{\bm{f}^{(i)}\}^{N_{\text{human}}}_{i=1}$ and body shape parameters $\{\bm{\beta}^{(i)}\}^{N_{\text{human}}}_{i=1}$ and corresponding SMPLx depth $\bm{d}_{\text{SMPLx}}$ following the processes in Sec. 3.1 and Sec. 3.2.
We then obtain residual features $\{R^{(i)}_{\text{face}}\}^{N_{\text{human}}}_{i=1}$ generated from the shared IdentityNet [71], $\mathcal{E}^{\text{ID}}_{\omega}$, given face embeddings $\{\bm{f}^{(i)}\}^{N_{\text{human}}}_{i=1}$ and face landmarks $\{\bm{p}_{\text{face}}^{(i)}\}^{N_{\text{human}}}_{i=1}$. Landmarks can be extracted from reference images or provided as additional 3D pose inputs. Also the residual feature $R_{\text{body}}$ is generated with SCNet $\mathcal{E}^{\text{SC}}_{\phi}$, given text $y$ and $\bm{d}_{\text{SMPLx}}$ derived from ${\{I_{\text{ref}}^{(i)}\}}^{N_{\text{human}}}_{i=1}$ and $\{\bm{p}^{(i)}\}^{N_{\text{human}}}_{i=1}$.
$\vspace{-3pt}R^{(i)}_{\text{face}}=\mathcal{E}^{\text{ID}}_{\omega}(\bm{f}^{(i)},\bm{p}_{\text{face}}^{(i)}),\ \ R_{\text{body}}=\mathcal{E}^{\text{SC}}_{\phi}(y,\bm{d}_{\text{SMPLx}})$ (2)
Let $F^{(k)}_{\theta}(\cdot)$ represent the $k$-th neural block and $\bm{s}^{(k)}$ the $k$-th input feature map. We obtain $\bm{s}^{(k+1)}$ by adding these residual features to $F^{(k)}_{\theta}(\bm{s}^{(k)})$, scaled by their respective conditioning weights $\alpha_{\text{face}}$ and $\alpha_{\text{body}}$, and modulated by face masks $M^{(i)}$:
$\vspace{-3pt}\bm{s}^{(k+1)}=F^{(k)}_{\theta}(\bm{s}^{(k)})+\alpha_{\text{face}}\sum\limits_{i=1}^{N_{\text{human}}}M^{(i)}\ast R^{(i)}_{\text{face}}+\alpha_{\text{body}}R_{\text{body}}$ (3)
Unlike two-stage methods (e.g., OMG [43]) or test-time optimization approaches (e.g., DreamBooth [60]), our method is more efficient and preserves the versatility of foundation models like SDXL. This enables training-free multi-human personalization from single reference images, with scalability to accommodate numerous identities without performance degradation.
3.4 User-Defined Body Shape Control
The SMPLx [57] shape parameters $\bm{\beta}$ are multiplied elementwise with the orthonormal shape displacement components, $\mathbf{S}$, and then added to the base shape, $\bm{v}_{\text{template}}$. Given $\bm{\beta}\in\mathbb{R}^{10}$, $\mathbf{S}\in\mathbb{R}^{N_{\text{vertices}}\times 3\times 10}$, and $\bm{v}_{\text{template}}\in\mathbb{R}^{N_{\text{vertices}}\times 3}$, the shaped vertices are defined as.:
$\vspace{-3pt}\bm{v}_{\text{shaped}}=\bm{v}_{\text{template}}+\sum_{i=1}^{10}\beta_{i}\cdot\mathbf{S}[:,:,i].$ (4)
Figure 6: Additional qualitative examples of PersonaCraft for single and multi-human personalization.
Table 1: Quantitative evaluation of PersonaCraft in face identity and body shape preservation, showing superior performance over baselines. CLIP similarity is comparable, with the result of text correspondence in user study (Tab. S1) addressing CLIP’s limitations in complex scenes.
Face ID preservation↑ Body shape preservation↑ CLIP similarity↑ # of idenities 1 2 3 4 5 Total 1 2 3 4 5 Total InstantID + OMG 0.369 0.274 0.200 0.169 0.155 0.194 0.505 0.406 0.349 0.307 0.302 0.370 0.267 IP-Adapter + OMG 0.227 0.127 0.134 0.118 0.134 0.125 0.538 0.442 0.377 0.340 0.335 0.403 0.266 IPA-Face + OMG 0.310 0.222 0.160 0.145 0.136 0.160 0.479 0.404 0.345 0.308 0.312 0.367 0.267 Ours 0.404 0.357 0.317 0.277 0.267 0.305 0.592 0.575 0.510 0.502 0.481 0.531 0.252
Table 2: Top-1 preference rates from user study on naturalness, face identity, body shape preservation, and text-image correspondence.
Top 1 (%) Natural.↑ Face ID↑ Body shape↑ Text corr.↑ DreamBooth 13.01 11.38 11.15 15.21 Textual Inversion 9.41 9.64 9.99 13.36 InstantID + OMG 5.92 9.99 8.83 8.36 IP-Adapter + OMG 15.45 13.24 14.29 12.66 IPA-Face + OMG 10.57 12.78 11.85 9.18 Ours 45.64 42.97 43.90 41.23
PersonaCraft offers user-defined body shape control, allowing adjustments based on user preferences. Specifically, we provide two types of controls: 1) Reference-based body shape control: This approach derives the body shape parameter $\bm{\beta}_{\text{target}}$ from the reference image using SMPLx fitting and directly applies it. 2) Interpolation and extrapolation-based body shape control: Here, we acquire two reference body shape parameters, $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$, and generate a controlled shape by interpolating or extrapolating between them. This target shape is defined as $\bm{\beta}_{\text{target}}=(1-\gamma)\bm{\beta}_{1}+\gamma\bm{\beta}_{2}$, where $\gamma\in\mathbb{R}$ adjusts the blend between $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$. The resulting $\bm{\beta}_{\text{target}}$ is then applied to Eq. 4 to obtain the target body shape. Results are shown in Fig. 7.
4 Experiments
4.1 Experimental Details
Implementation details. We implemented PersonaCraft using Stable Diffusion XL (SDXL) [59] as the base model. SCNet was trained on a curated dataset with a learning rate of 1e-5, using the Adam optimizer and a batch size of 2. The training process ran for 50,000 iterations on a single NVIDIA A100 GPU. For all experiments, we utilized 50 denoising steps.
Baselines. To evaluate the performance of PersonaCraft, we compared it against several baseline methods for single-shot, multi-identity, and pose-controllable image synthesis. All baselines were implemented using SDXL [59] to ensure a fair comparison. Key baselines include OMG [43], a state-of-the-art multi-concept personalization method, combined with InstantID/IP-Adapter/IP-Adapter-Face (IPA-Face) [71, 80], a leading single-shot personalization model and 2D pose ControlNet [81]. Optimization-based methods such as DreamBooth [60] and Texture Inversion [23] were also included for comparison. Additional results with other baselines, which may not align perfectly with the above categories, are also provided in the supplementary.
Test dataset. For testing, we evaluated our method using the COCO-WholeBody dataset [37], which contains full-body images of multiple people. The dataset includes images of 1 to 5 people, with a total of 1,000 images selected (200 images for each number of people). Text prompts were extracted from these images using BLIP [47] to generate corresponding textual descriptions. For pose conditioning, MultiHMR [12] was used to obtain 2D poses and 3D poses.
Metrics. Face identity preservation was measured for 1-5 identities following FastComposer [75], using FaceNet [63] for identity similarity within the face mask. Body shape preservation was evaluated using cosine similarity between SMPLx $\bm{\beta}$ parameters. Text-image correspondence was measured with CLIP-L/14 for image-text alignment. Additionally, a user study was conducted to assess perceptual quality.
More experimental details are in the supplementary.
Figure 7: Result of PersonaCraft’s user-defined body shape control. (a) Reference-based body shape control. (b) Interpolation and extrapolation-based control, where $\gamma$ controls the interpolation or extrapolation between the reference body shapes.
Figure 8: Results of combining PersonaCraft with various style LoRAs, showcasing adaptability to styles like Pastel, JoJo, and Pokémon Trainer. Some styles alter facial and body identities due to their bias, while producing visually impressive outcomes.
4.2 Results
Qualitative results. As shown in Fig. 45, and 6, our proposed method, PersonaCraft, demonstrates significant advantages over existing approaches, particularly in handling multi-identity images with occlusions and challenging poses. By leveraging a 3D-human model-conditioned ControlNet, PersonaCraft accurately preserves body shapes, enhances pose fidelity, and maintains body feature consistency with fewer artifacts in complex scenarios.
In contrast, methods like InstantID+OMG, IP-Adapter+OMG, and IPA-Face+OMG [71, 80, 43], which rely on 2D skeleton-based pose conditioning, exhibit severe anatomical inaccuracies in occluded or intricate interactions (highlighted by yellow arrows). These limitations arise from the inherent weaknesses of 2D pose representations, which struggle with overlapping body parts and complex scenarios. Additionally, blue arrows highlight cases where PersonaCraft maintains accurate body structure and pose fidelity, outperforming these methods in both identity preservation and body shape consistency.
DreamBooth [60] and Textual Inversion [23], which lack explicit pose guidance, often suffer from severe anatomical distortions, omission of individuals in multi-person scenes, and issue with clothing-body shape displacement, where clothing styles fail to adapt to individual body shapes.
Unlike 2D pose ControlNet, which struggles to distinguish occluded regions and handle challenging poses, PersonaCraft addresses these limitations effectively. It demonstrates robust performance as the number of subjects increases, preserving identity and body structure in multi-subject scenes. These findings highlight PersonaCraft’s robustness and versatility, establishing it as a state-of-the-art solution for personalized image generation in complex, real-world scenarios.
Quantitative results. We evaluate PersonaCraft’s performance in terms of face identity preservation, body shape preservation, and text-image correspondence. As stated in Tab. 1, the results show that our method consistently outperforms baselines in both face identity and body shape preservation across varying numbers of identities. Notably, it demonstrates robust body shape preservation with minimal variation across different body shapes. While text-image correspondence scores are comparable, it is important to note, as highlighted in previous literature [45, 42], that CLIP scores are less effective in evaluating images with complex scenes. To supplement this, we also conducted user studies, where PersonaCraft’s performance was highly rated.
User study. We conducted a user study to assess the naturalness, face identity preservation, body shape preservation, and text-image correspondence of images generated by three baseline methods (one from each group) and our method. Participants ranked their top three preferences for each attribute, yielding 10,332 responses from 123 participants. Results in Tab. S1 indicate that our method achieved the highest Top-1 preference across all metrics, with a significant margin over the baselines. Notably, the global CLIP score appeared misaligned with human judgment, underscoring its limitations in complex scenes. Further details of the user study setup are provided in the supplementary material.
Figure 9: Effectiveness of (a) SCNet compared to 2D Pose ControlNet [81] and (b) masking in multi-human image synthesis.
4.3 User-Defined Body Shape Control
A key feature of PersonaCraft is its user-defined body shape control, as demonstrated in Fig. 7. Users can select a reference image with the desired body shape, allowing the model to adapt these characteristics to generate highly personalized images. The system offers additional flexibility by enabling users to adjust body proportions through smooth interpolation or extrapolation between two different body references. This ability to fine-tune body shapes provides enhanced customization, making PersonaCraft a powerful and versatile tool for creating tailored image outputs with precise body shape control.
Table 3: Inference times for multi-identity personalized synthesis.
Method Total Time (secs) Text Inversion 1636.15 DreamBooth 770.71 InstantID + OMG 46.94 IP-Adapter + OMG 44.62 IPA-Face + OMG 35.46 PersonaCraft (ours) 17.25
Table 4: Evaluation of body shape preservation with and without the use of the body shape parameter.
Single Multi Total w/o body shape 0.559 0.493 0.506 w/ body shape 0.592 0.517 0.531
4.4 PersonaCraft with Stylization
The proposed method is a plug-and-play approach, making it compatible with various style-specific LoRAs. To evaluate its effectiveness, we conducted experiments combining PersonaCraft with diverse style LoRAs, including Crayon [55], Pastel [8], 3D Render [27], Pixel Art [53], Illustration [7], Frosting Lane [5], Pokémon Trainer [67], JoJo [54], Graphic Novel [15], and Cartoon [6]. The results, shown in Fig. 8, highlight the method’s ability to adapt to different styles effectively. Notably, styles such as Pastel, Illustration, JoJo, and Pokémon Trainer introduce changes in facial and body characteristics, occasionally altering perceived identity, due to their bias. Nevertheless, the outcomes remain visually compelling and demonstrate the versatility of our approach.
4.5 Efficiency Analysis
As analyzed in Tab. 3, we compare the inference times for multi-identity personalized synthesis across different methods, specifically for generating images with three distinct identities. Textual Inversion [23] and DreamBooth [60], which rely on optimization-based personalization, require a batch size of 4 and 500 optimization steps per identity. This process results in significantly longer inference times, making these methods highly inefficient for real-time applications.
On the other hand, methods based on OMG [43], which utilize a two-stage process to generate images, also require over twice the amount of time compared to our approach. In contrast, PersonaCraft performs inference efficiently, generating personalized images with substantially lower computation time while maintaining high-quality results. This stark difference highlights the efficiency advantage of our method, especially when generating multiple identities in a single synthesis process.
4.6 Ablation Studies
As shown in Tab. 4, the use of the body shape parameter enhances body shape preservation during personalization. Additionally, Fig. 9(a) demonstrates that our SCNet outperforms others in pose-guided image generation, effectively preserving human anatomy even in complex poses and occlusions. Moreover, Fig. 9(b) shows that our masking method improves performance, while adding features without masking increases feature variance and results in artifacts.
Additional results, including qualitative results, stylized artwork, PersonaCraft with LoRA and other face control modules, and ablation studies, can be found in the supplementary material.
5 Discussion
While our method is versatile and can be applied to other ControlNet models, the performance of our face personalization depends significantly on the underlying face network. Additionally, the accuracy of 3D human model fitting is contingent upon the performance of the fitting algorithm used. Variations in the quality of the fitting process may impact the final output, especially in cases where the reference data is incomplete or inaccurate. It is crucial to emphasize that PersonaCraft must be used responsibly, as the misuse of such technology for malicious purposes, including unauthorized identity manipulation, poses ethical and legal concerns.
6 Conclusion
In this paper, we presented PersonaCraft, a powerful framework for generating highly personalized images of multiple individuals, even in complex occlusion scenarios. By integrating a 3D human modeling approach with SMPLx-ControlNet, PersonaCraft effectively handles occlusions and ensures precise full-body personalization—advancing beyond previous methods that primarily focused on face identity or struggled with overlapping poses. The framework also introduces a novel feature for user-defined body shape control, which allows for more tailored and flexible customization, making it applicable to a wide range of scenarios.
Through comprehensive experimentation, PersonaCraft demonstrated superior performance in preserving individual identities, handling intricate poses, and maintaining body shape consistency across multiple individuals in a scene. We believe PersonaCraft holds significant promise for enhancing personalized content creation in diverse domains, including virtual environments, social media, and more. Future work will focus on refining the system to handle more complex interactions and real-world applications for dynamic scene personalization.
Supplementary Material
Appendix A Additional Results
A.1 Additional Qualitative Comparison
As shown in Fig. S1, our method significantly outperforms existing approaches. InstantID+OMG[71, 43] and IP-Adapter+OMG [80, 43], which use 2D skeleton-based pose conditioning, show severe anatomical inaccuracies in challenging cases (yellow arrows) due to the limitations of 2D pose representations. Blue arrows highlight that our method preserves body structure and pose fidelity while excelling in face identity and body shape consistency.
DreamBooth [60], lacking pose guidance, exhibits worse distortions, omitting individuals in multi-person scenes and misaligning clothing with body shapes.
These results highlight the robustness of PersonaCraft as a state-of-the-art solution for personalized image generation in complex scenarios.
A.2 Comparison with Additional Baselines
We compared PersonaCraft against additional baselines, including UniPortrait [32], MS-Diffusion [73], and FastComposer [76]. While these methods share similar capabilities, they are not fully suited for our benchmark, making direct comparisons challenging.
As shown in Fig. S2 and S3, yellow arrows highlight anatomical inconsistencies in complex poses and occluded scenarios due to reliance on 2D pose representations or the absence of pose control. PersonaCraft, in contrast, generates anatomically accurate and natural images under these conditions.
Blue arrows mark individuals used to evaluate body shape preservation, where PersonaCraft outperforms baselines, which frequently fail to maintain accurate body shapes.
Pink arrows indicate the inability of UniPortrait and FastComposer to handle multiple identities effectively, often mixing or duplicating identities. PersonaCraft successfully preserves distinct face identities for all individuals.
Additionally, MS-Diffusion copies clothing directly from full-body references without proper displacement. PersonaCraft integrates personalized body shapes and clothing displacement, maintaining consistency and realism.
These results highlight PersonaCraft’s superiority in generating accurate, identity-consistent images and handling occlusions and diverse poses with exceptional naturalness and customization.
A.3 Versatility of SCNet
To demonstrate the versatility of SCNet, we present results combining SCNet with various face identity personalization models, including InstantID [71], PhotoMaker V2 [48], and IPAdapter-Face [80]. As shown in Fig. S4, SCNet enables robust body shape personalization and pose control when paired with these face models, achieving comprehensive full-body personalization and user-defined body shape adjustments. Notably, face personalization varies slightly depending on the chosen face module.
A.4 Ablation Study on Conditioning Scale
We analyze the effect of the conditioning scales of IdentityNet and SCNet on identity preservation when provided with face references and reference body shapes (SMPLx depth). As shown in Fig. S5, when the conditioning scale is set to 0 for both modules, the generated face and body shapes differ significantly from the reference. This indicates insufficient guidance from the reference inputs.
As the conditioning scales for IdentityNet and SCNet increase, the generated images progressively resemble the reference face and body shape. This improvement demonstrates the critical role of conditioning strength in aligning the generated outputs with the given references. Optimal conditioning scales enable PersonaCraft to faithfully preserve both facial and body shape identities, ensuring high-quality personalization and consistency.
A.5 Additional Results of User Study
We present the cumulative percentage of the Top 3 results in Tab. S1, with details on the user study settings provided in the supplementary Sec. B.4. Consistent with the Top 1 results in Tab. 2 of the main paper, our method is preferred over other baselines in terms of all attributes.
Table S1: The cumulative percentage of Top 3 from user study on naturalness, face identity, body shape preservation, and text-image correspondence.
[…]
(Continued in Sec. B)
Figure S1: Additional comparison of generated images. Yellow arrows highlight anatomical issues due to 2D pose limitations. Blue arrows refer to the individuals evaluated for correct body shape preservation. PersonaCraft excels in identity, body shape consistency, and naturalness while being over twice as fast as OMG-based methods [71, 80, 43].
Figure S2: Comparison of PersonaCraft with UniPortrait [32], MS-Diffusion [73], and FastComposer [76]. Yellow arrows show anatomical inconsistencies in poses and occlusions. Pink arrows highlight identity mixing or duplication in baselines. MS-Diffusion copies clothing without proper adjustment.PersonaCraft excels in face identity, body shape consistency, and naturalness compared to the baselines.
Figure S3: Comparison of PersonaCraft with UniPortrait [32], MS-Diffusion [73], and FastComposer [76]. Yellow arrows show anatomical inconsistencies in poses and occlusions. Pink arrows highlight identity mixing or duplication in baselines. Blue arrows refer to the people for checking correct body shape preservation. MS-Diffusion copies clothing without proper adjustment.PersonaCraft excels in face identity, body shape consistency, and naturalness compared to the baselines.
Figure S4: Integration of SCNet with face personalization models like InstantID [71], PhotoMaker V2 [48] and IPAdapter-Face [80] achieves robust full-body customization, with slight variations by face module.
Figure S5: Ablation study on the conditioning scales of IdentityNet and SCNet, demonstrating improved identity preservation for face and body shape as the conditioning scales increase.
Appendix B Additional Experimental details
B.1 Additional Implementation Details
Face and body identity extraction. We selected MultiHMR [12] as our SMPLx fitting method. MultiHMR is a single-shot model that recovers 3D human meshes from a single RGB image, leveraging the SMPLx parametric model to predict full-body meshes, including hands and facial expressions, with 3D localization in the camera coordinate system. The body shape parameters, $\bm{\beta}$, are 10-dimensional vectors multiplied element-wise with orthonormal shape displacement components.
For facial identity processing, we use the antelopev2 facial detection and recognition models from InsightFace [1] to detect and extract 512-dimensional face identity embeddings, $\bm{f}$, from human images.
3D-aware pose conditioning using SMPLx-ControlNet. We base our SCNet on controlnet-union-sdxl-1.0 [77] and fine-tune it for SMPLx [57] depth conditioning. The architecture supports more than 10 control types for high-resolution text-to-image generation, with depth selected as the control type in our implementation.
We utilize 3D poses represented by SMPLx parameters, which include 55 joints (22 body, 1 jaw, 2 eyes, and 30 hands) along with camera parameters (intrinsic and extrinsic). These parameters generate a vertex- and face-based mesh that we render as SMPLx depth maps.
Text captions for MPII-SMPLx were derived using GPT-4 [3] by Khan et al. [38], providing detailed descriptions of samples from the MPII dataset.
Multiple human personalized image synthesis. We used the YamerMIX-v8 version of SDXL to enhance the visual quality of human-centric scenes. We selected IdentityNet from InstantID [71] as our face identity personalization module, which enables instant, zero-shot, identity-preserving image generation. The IdentityNet design imposes strong semantic and weak spatial conditions, integrating facial and landmark images with textual prompts to guide the image generation. For face landmarks, following InstantID, we used five key facial points (two for the eyes, one for the nose, and two for the mouth) as spatial control signals, applying a more generalized constraint compared to detailed key points.
B.2 Details on Metrics
Face identity preservation was measured for $1\sim 5$ identities following FastComposer [75], using FaceNet [63] for identity similarity within the face mask. The identity similarity score is computed by averaging the non-negative cosine similarity over both the number of humans and the total number of images:
$\mathcal{S}_{\text{face}}=\frac{1}{N_{\text{image}}}\sum_{i=1}^{N_{\text{image}}}\frac{1}{N_{\text{human},i}}\sum_{j=1}^{N_{\text{human},i}}\max\left(0,\frac{\bm{f}_{\text{ref},i}^{(j)}\cdot\bm{f}_{\text{gen},i}^{(j)}}{\|\bm{f}_{\text{ref},i}^{(j)}\|\|\bm{f}_{\text{gen},i}^{(j)}\|}\right)$ (S1)
where $\bm{f}_{\text{ref},i}^{(j)}$ and $\bm{f}_{\text{gen},i}^{(j)}$ are the face embeddings for the $j$-th reference and generated identity in the $i$-th image, respectively. $N_{\text{image}}$ is the total number of images, and $N_{\text{human},i}$ is the number of humans in the $i$-th image.
Body shape preservation was evaluated using cosine similarity between the SMPLx body shape parameters $\bm{\beta}$ from the reference and generated instances. The score is averaged over both the number of humans and the total number of test images:
$\mathcal{S}_{\text{body}}=\frac{1}{N_{\text{image}}}\sum_{i=1}^{N_{\text{image}}}\frac{1}{N_{\text{human},i}}\sum_{j=1}^{N_{\text{human},i}}\frac{\bm{\beta}_{\text{ref},i}^{(j)}\cdot\bm{\beta}_{\text{gen},i}^{(j)}}{\|\bm{\beta}_{\text{ref},i}^{(j)}\|\|\bm{\beta}_{\text{gen},i}^{(j)}\|}$ (S2)
where $\bm{\beta}_{\text{ref},i}^{(j)}$ and $\bm{\beta}_{\text{gen},i}^{(j)}$ are the body shape parameters for the $j$-th reference and generated instance in the $i$-th image, respectively. $N_{\text{image}}$ is the total number of images, and $N_{\text{human},i}$ is the number of humans in the $i$-th image.
CLIP similarity was measured using the CLIP-L/14 model for image-text alignment. Cosine similarity was used to evaluate the alignment between the generated image and the textual description. The CLIP encoders $\mathcal{E}_{\text{image}}$ and $\mathcal{E}_{\text{text}}$ were used for the image and text embeddings, respectively. The alignment score is averaged over all test images:
$\mathcal{S}_{\text{CLIP}}=\frac{1}{N_{\text{image}}}\sum_{i=1}^{N_{\text{image}}}\frac{\mathcal{E}_{\text{image}}(\bm{I}_{\text{gen},i})\cdot\mathcal{E}_{\text{text}}(y_{\text{ref},i})}{\|\mathcal{E}_{\text{image}}(\bm{I}_{\text{gen},i})\|\|\mathcal{E}_{\text{text}}(y_{\text{ref},i})\|}$ (S3)
where $\mathcal{E}_{\text{image}}(\bm{I}_{\text{gen},i})$ is the generated image embedding for the $i$-th image, and $\mathcal{E}_{\text{text}}(y_{\text{ref},i})$ is the reference text embedding.
Figure S6: Example images from the user study.
B.3 Details on Baselines
To evaluate PersonaCraft, we compared it with several baselines for single-shot, multi-identity, and pose-controllable image synthesis, all implemented using SDXL [59]. Key baselines include OMG [43] for multi-concept personalization, InstantID/IP-Adapter [71, 80] for single-shot personalization, 2D pose ControlNet [81], and optimization-based methods like DreamBooth [60] and Texture Inversion [23].
OMG + InstantID/IP-Adapter/IPA-Face. OMG [43] introduces a two-stage sampling solution for multi-concept personalization. The first stage handles layout generation and occlusion management, while the second stage integrates concepts using visual comprehension and noise blending. OMG can also be combined with single-concept models like InstantID without additional tuning. For OMG+InstantID, we follow the official inference code from the InstantID repository [44]. For OMG+InstantID and OMG+IP-Adapter/IPA-Face, we replace InstantID with IP-Adapter and IPA-Face, respectively, to adapt the framework for different face identity modules.
Textual Inversion. In original Textual Inversion [23], text embeddings are optimized for user-provided visual concepts, linking them to new pseudo-words that can be seamlessly incorporated into future prompts, effectively performing an inversion into the text-embedding space. To enable single-reference, multi-concept personalization, we optimize a unique text embedding $\mathcal{V}^{(i)}$ for each concept derived from a single reference image. These embeddings are paired with unique identifiers, allowing for the dynamic integration of multiple concepts into prompts during inference, facilitating multi-concept personalization.
DreamBooth. In original DreamBooth [60], the model is fine-tuned with images and text prompts using a unique identifier. A prior preservation loss is applied to encourage class diversity. For single-reference, multi-concept personalization, we adopt DreamBooth-LoRA [60, 35], where each reference image is associated with a unique $\mathcal{M}^{(i)}$ and identifier $\mathcal{V}^{(i)}$. These are fine-tuned based on the DreamBooth framework. During inference, both $\mathcal{M}$ and identifiers are used simultaneously, enabling personalized, concept-specific image generation from a single reference.
B.4 Details on User study
We conducted a user study to assess the naturalness, face identity preservation, body shape preservation, and text-image correspondence of images generated by three baseline methods (one from each group) and our method. Participants ranked the top three methods based on the following criteria: 1) Naturalness: Rank the images in order of the most natural-looking, considering factors such as physically impossible appearances, illogical features, inconsistencies, or lack of real-world physics and connections. 2) Face Identity Preservation: Rank the images in order of how well they reflect the face identity of the personalized character. 3) Body Shape Personalization: Rank the images based on how accurately they reflect the personalized character’s body shape. 4) Text Correspondence: Rank the images based on how closely they align with the textual description. The study collected a total of 10,332 responses from 123 participants across 7 cases, including both custom and COCO-Wholebody scenarios. We present illustrative example images from the user study in Fig. S6.
Acknowledgement
This work is supported by Samsung Electronics MX Division.